# Radar Perception in Autonomous Driving: Exploring Different Data Representations

### v1

@@ III-A3 Discussion
Another limitation is that ADC signals require a developer kit for data acquisition, which makes data processing challenging. Real-time object perception and vital sign monitoring pose computational challenges, requiring efficient hardware and software solutions. Moreover, neural networks processing ADC signals directly are suitable for classification tasks in closed environments, such as determining the presence or absence of objects and distinguishing between different types of objects [83]. However, in open environments, various distracting factors may lead to inaccurate object feature identification.
@@ III-B1 Datasets
  • 2D tensors: RADIATE [42], CRUW [85], FloW [86], MulRan [43], Oxford Radar RobotCar [44], SCORP [45];
@@
The 2D radar tensor dataset has the largest number of all radar tensor datasets. Specifically, RADIATE [42], CRUW [85], MulRan [43], Oxford Radar RobotCar [44] and SCORP [45] are presented in 2D range-azimuth coordinates, representing the BEV position of objects. The FloW [86] dataset is in 2D range-Doppler coordinates, showcasing the relationship between range and Doppler velocity for each object. CARRADA [40] is the first dataset that combines synchronized stereo RGB images and 3D radar RAD tensors. It provides annotations with bounding boxes, sparse points, and dense masks for both range-Doppler and range-azimuth representations. Regarding 4D radar tensors, K-Radar [50] appears to be the only available dataset, offering comprehensive information on range, Doppler, azimuth, and elevation.
@@ III-B2 Methods
2D tensor-based methods typically process 2D input radar data by considering range-Doppler and range-azimuth information. To conduct object detection on radar tensors, image-based network architectures (e.g., Faster R-CNN [87], ResNet [88], U-Net [89]) are applied or modified, as has been demonstrated in various studies using RA tensors [84, 90, 4] or RD tensors [91, 92, 93]. In general, 2D tensor-based methods employ a simplified radar data representation, reducing computational complexity. They can exploit correlations between two dimensions to improve detection and localization performance. However, these methods may only employ part of the available 3D spatial information from the radar. With limited representation, the accuracy of complex scenarios with multiple objects or occlusions may be reduced.
@@
3D tensor-based methods process radar data simultaneously using the range, Doppler, and azimuth dimensions, forming an RAD tensor. As such, this approach captures richer spatial information compared to 2D tensor-based methods, thus providing more valuable information to network architectures. Particular architectures have been developed to process aggregated views of 3D RAD tensors for object detection [3, 94, 95, 47, 4] and semantic segmentation [45, 96, 97]. Specifically, Major et al[3] first demonstrated the effectiveness of a deep learning-based object detection framework that operated on the RAD tensor and proved that the Doppler dimension helps increase detection performance. Similarly to the approach proposed by Major et al[3], Gao et al[94] decomposed the RAD tensor into three parts separately before combining them. The primary distinction between the two approaches is that the RA tensor in [94] consists of complex values that aid in recognizing objects using spatial patterns, thus increasing the accuracy of detection.
@@ III-B3 Discussion
Although radar tensors retain more comprehensive information about an object, they also reserve noise and clutter information, which may limit the ability to capture subtle object characteristics or distinguish between similar objects. Furthermore, the radar tensor provides a condensed representation of radar data, reducing the overall data size compared to raw ADC signals. Nevertheless, the radar tensor representation still requires considerable memory storage as well as large bandwidth, especially for the new 4D radar tensors.
@@ III-C1 Datasets
Conventional 3D radar sensors generate sparse point clouds, as observed in datasets such as nuScenes [51], Zender [41], SeeingThroughFog [53], HawkEye [100], AIODrive [54], RADIal [39], RadarScenes [55] and aiMotive [59]. Recently, 4D radar datasets are emerging, including Astyx [101], VoD [57], TJ4DRadSet [58], WaterScenes [60] and Dual-Radar [62]. While Astyx [101] stands as the first 4D radar point cloud dataset, it is limited by a small data size of only 500 frames. VoD [57] and TJ4DRadSet [58] datasets represent notable advancements in terms of data categories and data size. Meanwhile, these two datasets also incorporate simultaneous LiDAR data, facilitating comparative analysis between 4D radar point clouds and LiDAR point clouds. Interestingly, WaterScenes [60] is a 4D radar dataset focused on objects in waterway environments, which demonstrates the robustness of using 4D radar on water surfaces, particularly in adverse lighting and weather conditions. Dual-Radar [62] incorporates two types of 4D radars, expanding the comparison of different radar performances and further research on practical 4D radar perception algorithms.
@@ III-C2 Methods
Point cloud-based radar algorithms take radar point clouds as input and are widely applied for object detection [102, 103, 5, 104, 105], semantic segmentation [106, 102, 107, 108, 109, 20, 110, 111], tracking [112] and odometry [113, 114]. Based on various radar point cloud processing techniques, we categorize point-based methods into three groups: clustering-based, point-based, and voxel-based methods.
@@
Clustering-based methods [115, 116, 117] aim to group radar points into clusters using clustering algorithms (e.g., DBSCAN [37]), and then perform classification based on their spatial attributes. While clustering methods may be limited compared to deep learning-based methods, they are still effective in extracting individual objects from radar point clouds. Their simplicity also reduces memory requirements, making them widely used in radar-based segmentation tasks. However, clustering-based methods are sensitive to parameter settings, such as the density threshold in DBSCAN. Additionally, these methods may also struggle with cluttered, overlapping, and closely spaced objects.
@@
Point-based methods [106, 102, 107, 103, 5] process radar point clouds at the individual point level, extracting spatial features directly from these points. These methods generally draw inspiration from LiDAR-based algorithms, such as PointNet [118], PointNet++ [119] and Frustum PointNets [120]. These techniques leverage neural networks with convolutional filters to process radar points directly, considering spatial coordinates as integral features in the network structure. Point-based methods are good at accurate localization and classification of objects by considering individual radar points. However, they are susceptible to noise and occlusion, which may affect feature extraction accuracy.
@@ III-D Grid Map
Leveraging multi-frame point clouds, some researchers transform 3D radar reflections into a BEV pseudo-image, called the grid map [106]. There are two primary types of grid maps commonly used in radar-based perception: one is the occupancy-based grid map [128, 129, 130, 131, 132, 133, 134, 135, 136, 137], which represents the obstacles and free-space information; the other is the amplitude-based grid map [128, 129], which displays the reflected power values in that particular area. As shown in Fig. 8(b), an occupancy-based grid map has a clear outline, making it more feasible for outline-based detection tasks (e.g., parking spaces). An amplitude-based grid map, as indicated in Fig. 8(c), emphasizes the reflective characteristics of different objects, rendering it more appropriate for rural roads with a few distinct objects.
@@ III-D2 Methods
Occupancy-based methods [128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 139] utilize information such as the presence or absence of objects within a grid cell to make occupancy predictions. Traditionally, occupancy grid mapping is performed using an Inverse Sensor Model (ISM) and probabilistic methods (e.g., Bayesian filters [140]) to estimate occupancy probability [134, 141]. Additionally, EKF-based approaches model the environment as a Gaussian random field and estimate the state of each cell using a recursive Bayesian filter, taking into account both measurements and motion models [142]. Occupancy-based methods provide a concise environmental representation, simplifying obstacle detection and localization for efficient path planning [143]. However, occupancy-based methods may struggle to estimate occupancy states for cells with cluttered or overlapping radar returns [144]. Dynamic objects also pose problems due to their changing nature and intermittent radar measurements.
@@ III-E Micro-Doppler Signature
Micro-Doppler Signature (MDS) refers to the representation of micro-motion, such as rotation and vibration caused by object parts, resulting in a characteristic representation that differs from the Doppler frequency variation. MDS is generated by the Time-Frequency transform methods (e.g., Short Time Fourier Transform (STFT), Wavelet Transform (WT)) on Range-FFT results. As illustrated in Fig. 9, the spectrograms of walking and running pedestrians reveal different features. Specifically, the period of the micro-Doppler for a running pedestrian is shorter than that for a walking pedestrian. Moreover, the spectrograms between different types of objects (e.g., vehicles and bicycles) are unique, making them powerful features for object classification.
@@ III-E1 Datasets
The Open Radar Datasets [65] collects different types of outdoor moving targets, such as pedestrian walking and cycling, captured by a stationary radar system. The primary goal of this dataset is to utilize classification techniques for distinguishing between different motion activities. Apart from outdoor environments, some MDS-based datasets (e.g., CI4R [64], Dop-NET [63], MCD-Gesture [66]) for motion recognition (e.g., health monitoring, gesture recognition) are from indoor environments. These indoor datasets can assist in occupant behavior recognition to some extent, but still differ from behavior classification in autonomous driving due to inconsistent acquisition environments.
@@ III-E2 Methods
Over the past few years, significant research attention has been paid to the classification and recognition of objects’ postures and activities through MDS-based methods, particularly in the areas of object classification [84, 146], human activity and gait recognition [147, 148, 149, 150], and human gesture recognition [151]. Based on features of the MDS representation, we divide these methods into spectrogram-based methods and cadence velocity diagram-based methods.
@@
Spectrograms are a commonly used representation for analyzing micro-Doppler signatures. Spectrogram-based methods simultaneously analyze the MDS in both the time and frequency domains, which has been applied to human motion recognition [148, 150]. The time-varying gait information, including the speed of arm and leg swing, is effectively encoded within the spectrogram representation. Thus, spectrogram-based methods provide feature localization of MDS, allowing better identification and classification of different motions. However, spectrograms suffer from low time-frequency resolution trade-offs, leading to issues in capturing fast-changing Doppler signatures. Besides, spectrograms are also vulnerable to noise interference, which may obscure relevant information.
@@
Apart from the spectrogram presenting Doppler information in the time domain, another representation of MDS is the Cadence-Velocity Diagram (CVD) that exists in the frequency domain, which is derived by performing a Fourier transform along the time axis of the spectrogram [152, 153]. It provides a metric for capturing the frequency repetition patterns (“cadence frequencies”) over a given period. Thus, CVD-based methods provide a valuable measure for periodic changes in body parts at different velocities, such as walking or running [149, 11].
@@ III-E3 Discussion
MDS recognizes targets by observing minute movement characteristics of objects. This representation not only facilitates the differentiation of various object categories (e.g., pedestrians, bicycles, vehicles), but also enables the recognition of intricate object behaviors like gait and gesture recognition.
However, the transformation to the frequency domain via the Time-Frequency transform suffers from overlapping target components, resulting in suboptimal feature extraction for downstream tasks [73]. Similar to the ADC signal representation, since MDS is insufficient in delineating an object’s spatial coordinates required for detection and segmentation tasks, it is commonly employed for classification tasks. Moreover, the MDS representation may have limited resolution in range estimation compared to other representations like the radar tensor. This limitation could impact the ability to accurately determine the exact distance of an object, potentially affecting collision avoidance algorithms.
@@ IV-A Which Data Representation Should be Chosen?
ADC Signal
Classification Object Detection Vital Sign Monitoring Contains all information from raw radar data High temporal resolution Fine-grained feature recognition Lack of semantic information Complexity in data processing Sensitive to noise
Radar Tensor
Classification Detection Segmentation Localization Provides spatial structure with range, azimuth and elevation Provides rich motion information about the object Retains interference information around the object Inconsistent resolution between range and azimuth
Point Cloud
Detection Segmentation Tracking Odometry Offers 3D spatial representation of the surrounding environment Light-weight representation for objects Filters weak reflected signals Sparse in describing the shape of objects Susceptible to noise and occlusion Loss of potential information after CFAR processing
Grid Map
Mapping Localization Odometry Detection Provides structured BEV representation of the surroundings Offers a concise representation focusing on presence and occupancy Difficulty in representing complex and dynamic scenes Suffers from noise, occlusions and resolution
Micro-Doppler Signature
Motion Classification Vital Sign Monitoring Contains information about object motion Enhances differentiation between objects Limited resolution in range estimation Vulnerable to noise and interference

### v2

@@ III-A3 Discussion
Another limitation is that ADC signals require a developer kit for data acquisition, which makes data processing challenging. Real-time object perception and vital sign monitoring pose computational challenges, requiring efficient hardware and software solutions. Moreover, neural networks processing ADC signals are suitable for classification tasks in closed environments, such as determining the presence or absence of objects and distinguishing between different types of objects [87]. However, in open environments, various distracting factors may lead to inaccurate object feature identification.
@@ III-B1 Datasets
  • 2D tensors: RADIATE [42], CRUW [46], FloW [89], MulRan [43], Oxford Radar RobotCar [44], SCORP [45], OORD [51];
@@
The 2D radar tensor dataset has the largest number of all radar tensor datasets. Specifically, RADIATE [42], CRUW [46], MulRan [43], Oxford Radar RobotCar [44] and SCORP [45] are presented in 2D range-azimuth coordinates, representing the Bird’s Eye View (BEV) position of objects. The FloW [89] dataset is in 2D range-Doppler coordinates, showcasing the relationship between range and Doppler velocity for each object. CARRADA [40] is the first dataset that combines synchronized stereo RGB images and 3D radar RAD tensors. It provides annotations with bounding boxes, sparse points, and dense masks for both range-Doppler and range-azimuth representations. Regarding 4D radar tensors, K-Radar [50] appears to be the only available dataset, offering comprehensive information on range, Doppler, azimuth, and elevation.
@@ III-B2 Methods
2D tensor-based methods typically process 2D input radar data by considering range-Doppler and range-azimuth information. To conduct object classification or detection on radar tensors, image-based network architectures (e.g., Faster R-CNN [90], ResNet [91], U-Net [92]) are applied or modified, as has been demonstrated in various studies using RA tensors [88, 93, 4, 94, 95, 96] or RD tensors [97, 98, 99, 96]. For example, Gao et al[88] initiated the object detection process by defining a fixed-size bounding box for each detected object on the RA tensor. In addition, they extracted the radar data within the bounding box from multiple RA tensors and arranged this data in the form of radar data cubes. With the temporal information and object movement pattern, the data cubes serve as features for classifying different objects, such as pedestrians, cars and cyclists. Dong et al[93] applied ResNet [91] architecture on RA tensors for object detection. They also introduced uncertainty estimation for oriented bounding box localization further to enhance the accuracy of the object detection process. In general, 2D tensor-based methods employ a simplified radar data representation, reducing computational complexity. They can exploit correlations between two dimensions to improve detection and localization performance. However, these methods may only employ part of the available 3D spatial information from the radar. With limited representation, the accuracy of complex scenarios with multiple objects or occlusions may be reduced.
@@
3D tensor-based methods process radar data simultaneously using the range, Doppler, and azimuth dimensions, forming a 3D RAD tensor. As such, this approach captures richer spatial information compared to 2D tensor-based methods, thus providing more valuable information to network architectures. Particular architectures have been developed to process aggregated views of 3D RAD tensors for object detection [3, 100, 101, 47, 4] and semantic segmentation [45, 102, 103, 104]. Specifically, Major et al[3] first demonstrated the effectiveness of a deep learning-based object detection framework that operated on the RAD tensor and proved that the Doppler dimension helps increase detection performance. Similarly to the approach proposed by Major et al[3], Gao et al[100] decomposed the RAD tensor into three parts separately before combining them. The primary distinction between the two approaches is that the RA tensor in [100] consists of complex values that aid in recognizing objects using spatial patterns, thus increasing the accuracy of detection.
@@
Compared to using 2D tensor, 3D tensor-based and 4D tensor-based methods offer a more comprehensive representation of radar data, leveraging the complete spatial information and allowing for precise object detection, segmentation, and localization. These methods can effectively handle complex scenarios involving multiple targets, partial occlusions, and varying azimuth angles. However, the increased dimensions of the tensor may result in higher computational complexity and system delays.
@@ III-B3 Discussion
Although radar tensors retain more comprehensive information about an object, they also reserve noise and clutter information, which may limit the ability to capture subtle object characteristics or distinguish between similar objects. Furthermore, the radar tensor provides a condensed representation of radar data, reducing the overall data size compared to raw ADC signals. Nevertheless, the radar tensor representation still requires considerable memory storage as well as a large bandwidth, especially for the new 4D radar tensors.
@@ III-C1 Datasets
Conventional 3D radar sensors generate sparse point clouds, as observed in datasets such as nuScenes [52], Zender [41], SeeingThroughFog [54], HawkEye [107], AIODrive [55], RADIal [39], RadarScenes [56], aiMotive [60] and MiliPoint [64]. Recently, 4D radar datasets have emerged, including Astyx [108], VoD [58], TJ4DRadSet [59], WaterScenes [61] and Dual-Radar [63]. While Astyx [108] stands as the first 4D radar point cloud dataset, it is limited by a small data size of only 500 frames. VoD [58] and TJ4DRadSet [59] datasets represent notable advancements in terms of data categories and data size. Meanwhile, these two datasets also incorporate simultaneous LiDAR data, facilitating comparative analysis between 4D radar point clouds and LiDAR point clouds. Interestingly, WaterScenes [61] is a 4D radar dataset focused on objects in waterway environments, which demonstrates the robustness of using 4D radar on water surfaces, particularly in adverse lighting and weather conditions. Dual-Radar [63] incorporates two types of 4D radars, expanding the comparison of different radar performance and further research on practical 4D radar perception algorithms.
@@ III-C2 Methods
Point cloud-based radar algorithms take radar point clouds as input and are widely applied for object detection [109, 110, 5, 111, 112], semantic segmentation [113, 109, 114, 115, 116, 20, 117, 118], tracking [119, 120, 121], odometry [122, 123] and scene flow [124, 125, 126]. Based on various radar point cloud processing techniques, we categorize point cloud-based methods into three groups: clustering-based, point-based, and voxel-based methods.
@@
Clustering-based methods [127, 128, 129] aim to group radar points into clusters using clustering algorithms (e.g., DBSCAN [37]), and then perform classification based on their spatial attributes. While clustering methods may be limited compared to deep learning-based methods, they are still effective in extracting individual objects from radar point clouds. The simplicity of clustering-based methods also reduces memory requirements, making them widely used in radar-based segmentation tasks. However, clustering-based methods are sensitive to parameter settings, such as the density threshold in DBSCAN. Additionally, these methods may also struggle with cluttered, overlapping, and closely spaced objects.
@@
Point-based methods [113, 109, 114, 110, 5, 120, 121] process radar point clouds at the individual point level, extracting spatial features directly from these points. These methods generally draw inspiration from LiDAR-based algorithms, such as PointNet [130], PointNet++ [131] and Frustum PointNets [132]. These techniques leverage neural networks with convolutional filters to process radar points directly, considering spatial coordinates as integral features in the network structure. Schumann et al[113] proposed a structure based on PointNet++ for semantic segmentation on radar point clouds. Specifically, the Multi-Scale Grouping (MSG) module in PointNet++ helps group and generate features for a center point and its neighborhoods. They demonstrated that incorporating the RCS value and compensated Doppler velocity significantly improved classification accuracy. Inspired by [113], Danzer et al[109] adopted a two-stage method using PointNets [130, 132] for 2D car detection and segmentation. They considered each point as a proposal and adjusted the proposal size according to the object’s prior knowledge. Then, PointNet and Frustum PointNets are used to classify the proposals and each point in the proposal. Finally, the bounding box prediction is executed only for the proposals that are objects. Point-based methods are good at accurate localization and classification of objects by considering individual radar points. However, they are susceptible to noise and occlusion, which may affect feature extraction accuracy.
@@ III-D Grid Map
Leveraging multi-frame point clouds, some researchers transform 3D radar reflections into a BEV pseudo-image called the grid map [113]. There are two primary types of grid maps commonly used in radar-based perception: one is the occupancy-based grid map [140, 141, 142, 143, 144, 145, 146, 147, 148, 149], which represents the obstacles and free-space information; the other is the amplitude-based grid map [140, 141], which displays the reflected power values in that particular area. As shown in Fig. 9(b), an occupancy-based grid map has a clear outline, making it more feasible for outline-based detection tasks (e.g., parking spaces). An amplitude-based grid map, as indicated in Fig. 9(c), emphasizes the reflective characteristics of different objects, rendering it more appropriate for rural roads with a few distinct objects.
@@ III-D2 Methods
Occupancy-based methods [140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 151] utilize information such as the presence or absence of objects within a grid cell to make occupancy predictions. Traditionally, occupancy grid mapping is performed using an Inverse Sensor Model (ISM) and probabilistic methods (e.g., Bayesian filters [152]) to estimate occupancy probability [146, 153]. Additionally, Extended Kalman Filter (EKF) approaches take into account both measurement and motion models, modeling the environment as a Gaussian random field and estimating the state of each cell using a recursive Bayesian filter [154]. Occupancy-based methods provide a concise environmental representation, simplifying obstacle detection and localization for efficient path planning [155]. However, occupancy-based methods may struggle to estimate occupancy states for cells with cluttered or overlapping radar returns [156]. Dynamic objects also pose problems due to their changing nature and intermittent radar measurements.
@@ III-E Micro-Doppler Signature
Micro-Doppler signature refers to the representation of micro-motion, such as rotation and vibration caused by object parts, resulting in a characteristic representation that differs from the Doppler frequency variation. Micro-Doppler signature is generated by the Time-Frequency transform methods (e.g., STFT, WT) on Range-FFT results. As illustrated in Fig. 10, the spectrograms of walking and running pedestrians reveal different features. Specifically, the period of the micro-Doppler for a running pedestrian is shorter than that for a walking pedestrian. Moreover, the spectrograms between different types of objects (e.g., vehicles and bicycles) are unique, making them powerful features for object classification.
@@ III-E1 Datasets
The Open Radar Datasets [67] collects different types of outdoor moving targets, such as pedestrian walking and cycling, captured by a stationary radar system. The primary goal of this dataset is to utilize classification techniques for distinguishing between different motion activities. Apart from outdoor environments, some micro-Doppler signature-based datasets (e.g., CI4R [66], Dop-NET [65], MCD-Gesture [68]) for motion recognition (e.g., health monitoring, gesture recognition) are from indoor environments. These indoor datasets can assist in occupant behavior recognition to some extent, but still differ from behavior classification in autonomous driving due to inconsistent acquisition environments.
@@ III-E2 Methods
Over the past few years, significant research attention has been paid to the classification and recognition of objects’ postures and activities through micro-Doppler signature-based methods, particularly in the areas of object classification [88, 158], human activity and gait recognition [159, 160, 161, 162], and human gesture recognition [163]. Based on features of the micro-Doppler signature representation, we divide these methods into spectrogram-based methods and cadence velocity diagram-based methods.
@@
Spectrograms are a commonly used representation for analyzing micro-Doppler signatures. Spectrogram-based methods simultaneously analyze the micro-Doppler signature in both the time and frequency domains, which has been applied to human motion recognition [160, 162]. The time-varying gait information, including the speed of arm and leg swing, is effectively encoded within the spectrogram representation. Thus, spectrogram-based methods provide feature localization of micro-Doppler signatures, allowing better identification and classification of different motions. However, spectrograms suffer from low time-frequency resolution trade-offs, leading to issues in capturing fast-changing Doppler signatures. Besides, spectrograms are also vulnerable to noise interference, which may obscure relevant information.
@@
Apart from the spectrogram presenting Doppler information in the time domain, another representation of micro-Doppler signature is the Cadence-Velocity Diagram (CVD) that exists in the frequency domain, which is derived by performing a Fourier transform along the time axis of the spectrogram [164, 165]. It provides a metric for capturing the frequency repetition patterns (“cadence frequencies”) over a given period. Thus, CVD-based methods provide a valuable measure for periodic changes in body parts at different velocities, such as walking or running [161, 11]. However, CVD-based methods become complex when dealing with targets having multiple moving parts or irregular motion patterns. Fine details or rapid changes in motion can be challenging to capture accurately, limiting the ability to distinguish between closely spaced components or rapidly changing motions [166].
@@ III-E3 Discussion
Micro-Doppler signature recognizes targets by observing minute movement characteristics of objects. This representation not only facilitates the differentiation of various object categories (e.g., pedestrians, bicycles, vehicles), but also enables the recognition of intricate object behaviors like gait and gesture recognition.
However, the transformation to the frequency domain via the Time-Frequency transform suffers from overlapping target components, resulting in suboptimal feature extraction for downstream tasks [75]. Similarly to ADC signal representation, the micro-Doppler signature is insufficient in delineating an object’s spatial coordinates for detection and segmentation tasks, so it is commonly employed for classification tasks. Moreover, the micro-Doppler signature representation may have limited resolution in range estimation compared to other representations like the radar tensor. This limitation could impact the ability to accurately determine the exact distance of an object, potentially affecting collision avoidance algorithms.
@@ IV-A Which Data Representation Should be Chosen?
ADC Signal Classification Object Detection Vital Sign Monitoring Contains all information from raw radar data High temporal resolution Fine-grained feature recognition Lack of semantic information Complexity in data processing Sensitive to noise
Radar Tensor Classification Detection Segmentation Localization Provides spatial structure with range, azimuth and elevation Provides rich motion information about the object Retains interference information around the object Inconsistent resolution between range and azimuth
Point Cloud Detection Segmentation Tracking Odometry Offers 3D spatial representation of the surrounding environment Light-weight representation for objects Filters weak reflected signals Sparse in describing the shape of objects Susceptible to noise and occlusion Loss of potential information after CFAR processing
Grid Map Mapping Localization Odometry Detection Provides structured BEV representation of the surroundings Offers a concise representation focusing on presence and occupancy Difficulty in representing complex and dynamic scenes Suffers from noise, occlusions and resolution
Micro-Doppler Signature Motion Classification Vital Sign Monitoring Contains information about object motion Enhances differentiation between objects Limited resolution in range estimation Vulnerable to noise and interference